OpenStudy (anonymous):

There are 300 degrees in the sum of the interior angles of a polygon.

OpenStudy (unanimoose):

Well it depends on the polygon. Not all of them will equal 300.

sammixboo (sammixboo):

It is depending on what polygon Also there is no polygon that the angles add up to 300 Take a look here for more info http://www.mathsisfun.com/geometry/interior-angles-polygons.html

OpenStudy (anonymous):

Never. If it is a quadrilateral, the angles will add up to 360 degrees. For a pentagon, it's 540 degrees. For a hexagon, it's 720 degrees. In fact there is a formula for the degrees in a polygon. Where N = the number of sides, 180(n-2) = sum of angles. For a hexagon. 180 x (6-2) = 180 x 4 = 400+320 = 700

OpenStudy (anonymous):

So for the sum of angles to be 300, it would have to be 300 = 180(n-2) Or 300 = 180n-360 Balance it. 760 = 180n So N = 760 divided by 180, which is 4.2222 Obviously a shape cannot have 4.22222 sides, so we can say that a polygon will never have a sum of angles of 300 degrees.